Find the change in entropy when 1.25 kg of steam at 100°C condenses and cools to 50°C.

1
Expert's answer
2022-04-11T12:02:53-0400

"Q=m(L+c(100-50))=3.14~MJ,"

"∆S=\frac Q{∆T}=63~\frac{kJ}K."